# S1-Bench: A Simple Benchmark for Evaluating System 1 Thinking Capability of Large Reasoning Models

### v1

@@ C.1 Division of Different Format
Error Format. LRMs produce two types of undecodable errors: ID-300 indicates that the LRM outputs only a $\mathtt{ETM}$ without any additional content, while ID-301 represents cases where the LRM’s thinking process reaches the maximum length.
C.2 Evaluation of Different Format
In the classification shown in Table 6, ID-100 / 101 / 201 / 203 / 204 / 206 / 207 have decomposable final answers. If multiple $\mathtt{ETM}$s appear, only the last one is selected for splitting. In this case, the evaluator uses the final answer as the evaluation target.
ID-200 / 202 / 205 lack a final answer, in which case the evaluation target becomes the thinking process, as this part may also contain the correct solution to the question.
C.3 Statistics of Different Format
Table 8 presents the distribution of 12 format types under top-p sampling settings. Our analysis reveals three key insights: (1) The infinite thinking phenomenon is widespread across most models, particularly concentrated in LRMs with fewer than 32B parameters. This indicates that output formatting should be prioritized during training to reduce format errors, especially when working with less powerful foundation models. (2) The Nemotron and EXAONE families frequently produce correctly formatted responses without visible reasoning processes. This behavior can be viewed as a mechanism for mitigating over-thinking. However, the EXAONE family still exhibits substantial overthinking tendencies, suggesting that LRMs’ capability to respond without visible reasoning and their tendency to overthink may be orthogonal characteristics. (3) None of the evaluated LRMs exhibited behaviors classified as ID-205/206, possibly because special $\mathtt{ETM}$s are perceived by LRMs as the only alternative to standard $\mathtt{ETM}$s. Furthermore, ID-300 was not observed, indicating that high-temperature sampling generally yields semantically meaningful responses.
Table 9 shows the format statistics of greedy sampling. The phenomenon is generally similar to top-p sampling. L-R1-32B produced ID-300, possibly due to poor compatibility between the training method and low temperature.
Appendix D Experiment Details
Table 10 shows LRMs’ performance on S1-Bench under greedy sampling settings. While overall accuracy improves compared to top-p sampling settings, $\mathtt{ART}$ slightly decreases, yet overthinking and accuracy degradation issues persist. Figure 9 illustrates the $\mathtt{ART}$ heatmap across 28 subcategories and 26 models.
We compile the statistics of the prejudgment words extracted, as shown in Table 7.
Model Size acc (Loose) acc (Strict) $\mathtt{SUC}$-$\mathtt{L}\uparrow$ $\mathtt{SUC}$-$\mathtt{S}\uparrow$ $\mathtt{ART}$downarrow
QwQ-32B 32B 100.00 100.00 100.00 100.00 750.41
Hunyuan-T1 100.00 100.00 100.00 100.00 541.09
Sky-T1-32B 32B 99.53 99.05 99.76 99.29 157.12
DS-R1 671B 100.00 100.00 100.00 100.00 621.89
DS-R1-70B 70B 99.76 99.76 100.00 100.00 469.78
DS-R1-32B 32B 100.00 100.00 100.00 100.00 428.46
DS-R1-14B 14B 99.29 99.29 99.76 99.76 463.52
DS-R1-8B 8B 97.63 97.39 99.76 99.53 452.11
DS-R1-7B 7B 94.31 94.31 97.87 97.87 436.87
DS-R1-1.5B 1.5B 76.54 76.54 91.94 91.94 473.67
Nemotron-49B 49B 99.53 99.53 100.00 100.00 337.94
Nemotron-8B 8B 84.60 77.73 98.34 81.99 446.62
L-R1-32B 32B 92.18 85.78 94.55 85.78 996.36
L-R1-32B-DS 32B 99.29 99.29 99.29 99.29 528.45
L-R1-14B-DS 14B 98.82 98.82 98.82 98.82 664.28
L-R1-7B-DS 7B 92.65 92.65 98.82 98.82 514.60
s1.1-32B 32B 98.82 98.82 98.82 98.82 983.38
s1.1-14B 14B 95.97 95.50 96.45 95.97 786.30
s1.1-7B 7B 94.31 87.68 94.55 87.91 630.52
EXAONE-32B 32B 97.63 97.39 100.00 99.76 746.89
EXAONE-7.8B 7.8B 86.73 86.49 97.63 97.16 947.92
EXAONE-2.4B 2.4B 72.99 72.99 94.79 94.79 1394.72
Figure 9: $\mathtt{ART}$ on the 28 subcategories, which is the average result of five generations under top-p sampling with Loose Format setting.
Appendix E Prompts discussed in this paper
This section presents the prompts used in this paper. For S1-Bench construction-related prompts, please refer to Table 11; for experimental prompts, please refer to Table 12,13 and 14.
Prompt for construction workflow for S1-Bench
Initial Generation
Based on the category name and explanation, generate 50 pairs of questions and answers in both Chinese and English. Ensure that the generated questions vary in perspective and content, maintaining diversity. All questions must have a **uniquely determined, unambiguous answer with no potential conditions causing multiple possible answers**. Ensure that the questions are expressed clearly. Additionally, the difficulty level of the questions should be **as simple as possible—so easy that they can be answered correctly without any need for thought**.
# Category Name:
{category}
# Explanation:
{explanation}
# Cases:
## English question:
{question_en}
## English Answer:
{answer_en}
## Chinese question:
{question_zh}
## Chinese Answer:
{answer_zh}
Start generating:
Discriminating Generation Quality
1. Whether the question belongs to the specified category.
2. Whether the question is clear, unambiguous, has an absolutely unique answer, and has no potential alternative interpretations.
# Question Category and Explanation:
{category_with_exp}
# Question and Answer:
{question_with_answer}
# Other Questions:
{questions_list}
Reduce Difficulty Generation
- Maintain the original perspective and semantic information
- Preserve the question’s alignment with its category
# Question Category and Explanation:
{category_with_exp}
# Question and Answer:
{question_with_answer}
The new question and answer:
Evaluation on S1-Bench
**Question:**
{question}
**Model Answer:**
{model_answer}
**Ground Truth:**
{ground_truth}
Your task is to evaluate whether the model’s answer is correct.
Table 11: Prompt for construction workflow for S1-Bench. The “explanation” (and “category_with_exp”) includes the definition of each major category, simplicity requirements in Section 3.1.1, and the definition of subcategories.
Prompts for Solution Segmentation
# Task Description:
1. Segmentation positions:
2. Output Restriction:
# Some cases:
## Case 1:
### Question:
What is the value of x if x plus 3 equals 8?
### Ground Truth:
5
### Chain of Thought:
Original equation: x + 3 = 8
Substitute x = 5: 5 + 3 = 8
5 plus 3 is indeed 8, so that works out.
*Final Answer**
The value of x is 555.
### Output:
Original equation: x + 3 = 8
Substitute x = 5: 5 + 3 = 8
5 plus 3 is indeed 8, so that works out.<split>
*Final Answer**
The value of x is 555.<split>
## Case 2:
## Case 3:
# Eval Target:
## Question:
{question}
## Ground Truth:
{ground_truth}
## Chain of Thought:
{thinking_process}
## Output:
Table 12: Prompts for Solution Segmentation.
Prompts for Error Analysis
Types of Correct
# Task Description:
# Categories:
Output your evaluation in the following format:
# Question:
{question}
# Ground Truth:
{gound_truth}
# COT:
{thinking_process}
# TheReason:
Types of Error
# Task Description:
# Categories:
Output your evaluation in the following format:
# Question:
{question}
# Ground Truth:
{answer}
# COT:
{thinking_part}
# TheReason:
Table 13: Prompts for Error Analysis.
Prompts for Extracting Prejudgment Sentences and Words
Extracting Sentences
# Given the beginning of a model’s thought process when answering a question, determine if there is any evaluation by the model regarding the difficulty level or category of the question.
# For example:
Example Evaluation 1: This seems to be an easy question.
Example Evaluation 2: {CJK}UTF8gkai这看起来似乎是个基础的问题
Example Evaluation 3: This question is a bit complex.
Example Evaluation 4: {CJK}UTF8gkai这个问题看起来复杂
# Beginning of thought process:
{thinking_process}
# Please output only the evaluative short phrase you find. If there is no evaluation, output "None".
# Output:
Extracting Words
# You are given a comment sentence, and your goal is to find the only comment-related word.
# For example:
## case 1:
Sentence: {CJK}UTF8gkai这是一个很简单的算术题
Word: {CJK}UTF8gkai简单
## case 2:
Sentence: that seems pretty straightforward
Word: straightforward
# The comment sentence:
{review_sentence}
Output the only comment-related word:
Table 14: Prompts for Extracting Prejudgment Sentences and Words.

### v3

@@ C.6 Gut Moment Analysis Details
Type Sentence Count
easy-zh {CJK}UTF8gkai 这个问题看起来挺简单的 308
easy-en that seems straightforward 36
difficult-zh {CJK}UTF8gkai 这个问题看起来有点复杂 308
difficult-en percentages can sometimes be tricky 7
neutral-zh {CJK}UTF8gkai 这个问题看起来好像不难 24
neutral-en Hmm, interesting 3
Table 6: The most common sentence in each type of “gut moment."
Appendix D Error Cases
This section presents several error cases observed in LRMs. See Tables 2021222324, and 25.
Paper Abbreviation MATH GSM8K AIME GPQA AMC MMLU Olympiad- SVAMP LiveCode- Common-
Bench Bench SenseQA
Codi Shen et al. (2025b)
CISC Taubenfeld et al. (2025)
CoT-Valve Ma et al. (2025b)
Dast Shen et al. (2025a)
ATM Chen et al. (2024a)
DEER Yang et al. (2025b)
DPTS Ding et al. (2025)
Dynasor Fu et al. (2024)
ESC Li et al. (2024)
Hawkeye She et al. (2025)
token complexity Lee et al. (2025)
INFTYTHINK Yan et al. (2025)
KIMI K1.5 Team et al. (2025a)
L1 Aggarwal and Welleck (2025)
LightThinker Zhang et al. (2025a)
LS-Mixture SFT Yu et al. (2025a)
DSC Wang et al. (2024)
O1-Pruner Luo et al. (2025)
MRT Qu et al. (2025b)
Self-Doubt Fu et al. (2025)
RASC Wan et al. (2024)
NoThinking Ma et al. (2025a)
Retro-Search Lu et al. (2025)
RSD Liao et al. (2025)
ST-BoN Wang et al. (2025b)
Elastic Reasoning Xu et al. (2025b)
FS-BoN Munkhbat et al. (2025)
SoT Aytes et al. (2025)
SpecReason Pan et al. (2025)
Speculative Thinking Yang et al. (2025d)
SPIRIT Cui et al. (2025)
ITC Analysis Wang et al. (2025a)
Think when needed Yang et al. (2025c)
THINKPRUNE Hou et al. (2025)
TALE Han et al. (2024)
TokenSkip Xia et al. (2025)
TOPS Yang et al. (2025e)
efficient reasoning Arora and Zanette (2025)
TWT Xu et al. (2025a)
Z1 Yu et al. (2025b)
Count 28 24 20 11 8 6 6 5 5 5
Table 7: A total of 40 studies on LRM efficiency before May 2025 were included. Benchmarks that appeared more than four times are listed.